2. The government is contemplating project A that yields the following stream of benefits and costs. Project A Year 2 S10,0000 0 Benefits0 $2,000 $4,000 $6,000 (a) Find the net present value and the benefit cost ratio for this project, assuming a discount rate of 5%. (b) Re-calculate the net present value assuming a discount rate of (i) 3%, (ii) 8 % and (iii) 10%. Use a graph to illustrate how the net present value of the project is affected by the choice of discount rate. (b) Suppose that project A has the stream of benefits and costs as above but the analyst forgot to include the horizon value of the project. Estimates suggest that the horizon value evaluated in year 3 is-$1000. Explain why the horizon value might be negative. Use this information to recalculate the net present value of the project (assuming a 5% discount rate as is done in part a). Is the project worthwhile? Explain.
Answer #1

Net present value (NPV) is the difference of present value of net benefits and the initial (year 0) cost.

Net benefit (NB) = Benefit – Cost

A project could be accepted, if it has positive NPV.

a.

NPV at 5% discount rate

Year

NB

5% discount factor = 1/(1 + 0.05)^n

NB × factor

0

0 – 10,000 = -10,000

1

-10,000

1

2,000

0.9524

1,904.80

2

4,000

0.9070

3,628.00

3

6,000 – 6,000 = 0

0.8638

0

4

8,000

0.8227

6,581.60

5

10,000 – 2,000 = 8,000

0.7835

6,268.00

NPV

8,382.40

Answer: NPV is $8,382.40.

Answer: The project is worthwhile, since it has positive NPV.

b.

NPV at 10% discount rate

Year

NB

5% discount factor = 1/(1 + 0.05)^n

NB × factor

0

0 – 10,000 = -10,000

1

-10,000

1

2,000

0.9091

1,818.20

2

4,000

0.8264

3,305.60

3

6,000 – 6,000 = 0

0.7513

0

4

8,000

0.6830

5,464.00

5

10,000 – 2,000 = 8,000

0.6209

4,967.20

NPV

5,555

Answer: NPV is $5,555.

Answer: The project is recommended, since it has positive NPV.

c.

If the benefit becomes ½, net benefits in each year would be as below:

Year

Benefits, B

Costs, C

Net benefits [B – C]

0

0

10,000

-10,000

1

1,000

0

1,000

2

2,000

0

2,000

3

3,000

6,000

-3,000

4

4,000

0

4,000

5

5,000

2,000

3,000

NPV at 5% discount rate

Year

NB

5% discount factor = 1/(1 + 0.05)^n

NB × factor

0

-10,000

1

-10,000

1

1,000

0.9524

952.40

2

2,000

0.9070

1,814.00

3

-3,000

0.8638

-2,591.40

4

4,000

0.8227

3,290.80

5

3,000

0.7835

2,350.50

NPV

-4,183.70

Answer: NPV is -$4,183.70.

Answer: The project is not recommended, since it has negative NPV.

Explanation: Once benefits in each year go down, NB becomes low. This is the reason why NPV becomes negative.

Part 2: NPV at perpetuity would be as below:

NPV = Net benefits / Discount rate

         = $100,000 / 0.08

         = $1,250,000 (Answer)